# Unknotting number is not additive under connected sum

### v1

@@ 1. Introduction
$u(7_{1}\#\overline{10_{139}})<u(7_{1})+u(10_{139})$ and $u(10_{139}\#\overline{10_{139}})<u(10_{139})+u(10_{139})$
We expect that there are many other knots whose unknotting numbers are known and that contain $7_{1}$ in a minimal unknotting sequence, which would then yield other connected sums that fail additivity. For example, any intermediate knot $K$ in a Gordian adjacency from the torus knot $T(p,q)$ to the knot $7_{1}$ also has $7_{1}$ Gordian adjacent to it, and so can be used as a summand in such examples. Several authors have explored the construction of minimal unknotting sequences for torus knots [3],[33], and the problem of Gordian adjacency for torus knots [11],[5]. Feller’s work [11] shows that all torus knots except (possibly) $T(2,3),T(2,5),T(3,4),T(4,5)$ and $T(5,6)$ have $7_{1}=T(2,7)$ Gordian adjacent to them; $T(2,3),T(2,5)$ and $T(3,4)$ cannot, because their unknotting numbers are too low. The cases $T(4,5)$ and $T(5,6)$ are unresolved.
@@ 3. How (and why) the main example was found
The process of finding the diagram for $7_{1}\#\overline{7_{1}}$ and the discovery of its unexpectedly low unknotting number had its genesis in another ongoing project of the authors: our search for an example to answer a different question, namely whether or not unknotting number is invariant under mutation. This is Problem 1.69(C) in Kirby’s problem list [16]. Our approach to that problem included building a collection of knots which would have unknotting number 4 (if unknotting number were additive), but which would have comparatively small signature. While the inequality $|\sigma(K)|/2\leq u(K)$ [21] is useful in bounding unknotting number from below, it can be difficult to find knots where the difference between these two quantities is known to be greater than 1 (which we needed for our search), although work of Stoimenow [34] is a significant outlier, which deserves to be more widely known. On the other hand, the facts that signature is additive under connected sum, and that the signature of the mirror image of a knot $K$ is the negative of the signature of $K$ (see, e.g., [21]), means that is it a routine matter to build connected sums where the unknotting number would be large, if unknotting number were additive, but where the signature is small. It was for this reason that we began to search for knots a few crossing changes away from connected sums of low crossing number knots with known unknotting number, in order to build a collection of new knots with (if unknotting number were additive) high enough unknotting number to be useful for our search.
@@
Our approach to finding the new knots was to adopt the process we had followed for two previous projects, where we settled the Bernhard-Jablan Conjecture (in the negative) [6] and where we completed the computation of the smooth 4-genus of the prime knots through 12 crossings [7]. We started with pairs of knots, one with known unknotting number 3 and the other with known unknotting number 4, took their connected sum, and then built diagrams for these sums for which crossing changes stood a good chance of building prime knots, with low enough (after simplification) crossing number that Snappy could recognize them (that is, crossing number 15 or less). We chose unknotting numbers 3 and 4 so that we were starting with knots that were likely to have unknotting number 7, and then randomly changed three crossings so that the resulting knots ought to have unknotting number (at least) 4. We chose to build the diagrams by starting with the standard diagrams for the connected sums, randomly performing some crossing-number-increasing Reidemeister moves using Snappy’s backtrack() command, and then converting them to braids using the braid_word() command. For this last step, Snappy uses Vogel’s algorithm [36] to convert the knot diagram into a closed braid. We chose this approach because (a) Snappy can quickly carry out both of these operations and (b) our experience with the previous two projects had shown us that working with braid representatives built in this way succeeds in building diagrams that are randomized enough to yield good results (that is, many different knot types) when we change crossings in the braid representations. Experience also showed that by changing 3 crossings, when starting from a connected sum in this way, resulted in a good proportion of the knots that were generated being prime; 2 changes was typically insufficient. Using braid words, changing crossings is very quick; it amounts to changing the signs of entries in the braid representation.
@@
Our work in [6] had provided computations of upper bounds on the unknotting number of all of the prime knots through 15 crossings (what we called, in that paper, their weak BJ-unknotting number, that is, their unknotting number if the Bernhard-Jablan conjecture (and additivity of unknotting number!) were true). We could have high confidence that the vast majority of these computations represented the actual unknotting number of the knots, based on comparison with known values for lower crossing number knots. We compared the knots coming from the randomized procedure to our list of ‘weak BJ-unknotting number 4’ knots to create a group of what we called ‘conditionally unknotting number four’ knots. Working with several million braids yielded 240 such 14- and 15-crossing knots. One of these was the knot $K14a18636$.
@@
We then began to reverse the search, using these 240 knots as initial knots to carry out crossing changes on their braid representatives, in the same manner in which the connected sum diagrams had been built. This is where things got interesting! For the mutation project, our goal was to show that one of a collection of target knots had unknotting number 3, by showing that it was crossing-adjacent to a knot with unknotting number 4. Prior such searches using knots with unknotting number known to be 4 had failed to make such a connection, and so our plan was to search among knots one crossing change away from these 240 knots with conditional unknotting number 4, instead. Connecting one of the 240 initial knots with one of these mutant ‘targets’ would have proved that either unknotting number is not mutation invariant or that the knot among our 240 that made the connection could not actually have unknotting number 4, and so, as argued above, unknotting number would not be additive under connected sum. (But, were such a thing to happen, we would not know which of the two possibilities was true!)
@@
At the start of this phase of the project it was not at all clear what it would produce. The most likely prospect seemed to be that we would never find a connection between our 240 initial knots and our list of target knots, if only because both conjectures might be true. Our expectation, if the project succeeded, was that it would be because the target knot we connected with had unknotting number 3, and so unknotting number would not be mutation invariant. It came as quite a surprise then when, instead, this reverse search found a crossing change from one of the 240 knots we had identified, namely $K14a18636$, to the knot $K15n81556$, which had a known, and unexpectedly low, weak BJ-unknotting number of 2.
@@
The first example we found started from a closed braid representation of the knot $7_{1}\#10_{139}$; three crossing changes, that our search had found, transformed this to a representative of the knot $K14a18636$. One of those three crossing changes transformed the braid into a braid representative for the knot $7_{1}\#\overline{7_{1}}$, and so this braid word contained two crossing changes to a representative of the knot $K14a18636$. Initially, the braid word had length 113, and so represented a 113-crossing diagram. But several simplifications of the braid word could, it turned out, be carried out. These mostly involved Reidmeister I moves ‘hidden’ by a sequence of Reidemeister II moves. Their discovery was aided by the program KnotPlot [32], to visualize long stretches of the braid. This enabled the authors to greatly reduce the number of crossings needed (from 113 to 20), and the number of strands in the braid needed (from 14 to 5). This resulted in the diagram and braid word used in our discussion in Section 2.
@@ 4. The road forward
With the additivity of unknotting number being settled in the negative, it might seem like the quest to build a more comprehensive understanding of unknotting number may have hit a bump in the road. Knowing for sure that unknotting number does not behave well with respect to connected sum means that the chances for a purely computational approach to determining unknotting numbers, as one could argue the other invariants mentioned in the introduction enjoy, seems even further out of reach. Further, we cannot rely on knowledge of unknotting numbers for prime knots to provide all of the information we need to compute the invariant more generally.
As the authors of the reinforcment learning approach [2] had postulated, additivity under connected sum presented a potentially powerful opportunity to compute the unknotting numbers of knots with relatively small crossing number but unknown unknotting numbers, by building ‘test’ knots, as connected sums, which effectively decouple the unknotting number from the signature (and, as it happens, the 4-genus), to change crossings in. With the loss of this technique, we need new tools to determine the unknotting numbers, for example, of the nine 10-crossing prime knots whose values we still do not know.
On the other hand, knowing that additivity of unknotting number can in fact fail makes it more interesting to discover when it does not fail. That is, establishing that $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for families of knots has increased importance, since we can no longer expect this to hold in general. There is also the question of how badly additivity fails; is it, perhaps, even possible that the unknotting number of a connected sum can be lower than that of its summand?
@@
Does $u(K\#K^{\prime})\geq\max\{u(K),u(K^{\prime})\}$ for all knots $K$ and $K^{\prime}$
The fact that we could bootstrap an initial, single, example of non-additivity to find infinitely may knots $K$ for which $u(K\#7_{1})<u(K)+u(7_{1})$ raises the question of whether or not every knot $K$ can be paired with a ‘poison’ knot $K^{\prime}$ so that $u(K\#K^{\prime})<u(K)+u(K^{\prime})$. So we ask:
@@
Does there exist a non-trivial knot $K$ for which $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for every knot $K^{\prime}$
While we now know that $u(7_{1}\#\overline{7_{1}})<u(7_{1})+u(\overline{7_{1}})$, and we can build infinitely many examples of this phenomenon from this, all of our examples currently rely on a sequence of $5$ crossing changes that unknot $7_{1}\#\overline{7_{1}}$ We would certainly like to understand why this sequence occurred! What, if anything, is special about the knot $7_{1}$ that might make this happen? The intermediate knot $K_{1}=K14a18636$, which played a central role in establishing the non-additivity of unknotting number, has, from our work in [6], a weak BJ-unknotting number of 4, yet the unknotting sequence above shows that it has unknotting number at most 3. So the knot $K_{1}$ is, in fact, a counterexample to the weak Bernhard-Jablan conjecture [6]. That is, no crossing change in its (unique) minimal crossing diagram lowers the weak BJ-unknotting number. This may point the way to building more examples, by aiming to find a BJ-unknotting counterexample in an unknotting sequence. We note that there are (at least) six other knots that the authors have identified with weak BJ-unknotting 4 and unknotting number $\leq 3$: they are the knots $K13n1669$, $K14a2644$, $K14a11300$, $K14a11565$, $K14a12383$, and $K14a13809$. Each is crossing-adjacent to a 15-crossing diagram for a knot $K$ with $u(K)\leq 2$ (the corresponding unnotting number (at most) 2 knots are, respectively, $K15n9318$, $K15n9318$, $K15n35674$, $K15n82797$, $K15n4866$, and $K15n129095$), and so each of these knots has unknotting number at most $3$. These could, like $K14a18636$, be useful targets for building other unexpectedly short unknotting sequences for connected sums.
Finally, we are left with the question:
@@
What is the actual unknotting number of $7_{1}\#\overline{7_{1}}$

### v2

@@ 1. Introduction
$u(7_{1}\#\overline{10_{139}})<u(7_{1})+u(10_{139})$ and $u(10_{139}\#\overline{10_{139}})<u(10_{139})+u(10_{139})$ .
We expect that there are many other knots whose unknotting numbers are known and that contain $7_{1}$ in a minimal unknotting sequence, which would then yield other connected sums that fail additivity. In particular, any intermediate knot $K$ in a Gordian adjacency from the torus knot $T(p,q)$ to the knot $7_{1}$ also has $7_{1}$ Gordian adjacent to it, and so can be used as a summand in such examples.
For further examples, by leveraging a result of Baader [6], we can build a great many non-torus knots for which $u(L_{1}\#\overline{L_{2}})<u(L_{1})+u(L_{2})$. Baader shows that for any two knots $K$, $L$ that are two crossing changes apart, there are infinitely many distinct knots $K_{i}$ that are intermediate to them, in the sense that they are each one crossing change away from both $K$ and $L$. By finding infinitely many such knots $K_{i}$ intermediate between the knots $T(2,2(n+1)+1)$ and T(2,2(n1)+1T(2,2(n-1)+1, for $n\geq 4$, we can be certain of two things: First, each of the knots $K_{i}$ has unknotting number precisely $n$, since they are each one crossing change away from $T(2,2(n+1)+1)$, with unknotting number $n+1$, and from T(2,2(n1)+1)T(2,2(n-1)+1), with unknotting number n1n-1. The value of $u(K_{i})$ must then be at most one away from both $n+1$ and n1n-1. Second, $7_{1}$ is Gordian adjacent to each $K_{i}$, since, in our standard unknotting sequence for $T(2,2(n+1)+1)$, passing through T(2,2(n1)+1T(2,2(n-1)+1 and $7_{1}$, we can replace the knot $T(2,2n+1)$ with $K_{i}$ and still have an unknotting sequence of the same length, which passes through $K_{i}$ and then $7_{1}$. This gives us Corollary 1.5. $\square$
Every diagram $D$ of a knot $K$ can be turned into a diagram of the unknot by changing at most $\lfloor c(D)/2\rfloor$ crossings, where $c(D)$ is the number of crossings in the diagram. This can be done, for example, by changing crossings to build an ascending diagram of the unknot from the diagram $D$. This leads to Ozawa’s result [33] that $u(K)\leq\lfloor(c(K)-1)/2\rfloor$, where $c(K)$ is the crossing number of $K$, the minimum number of crossings among all diagrams for $K$. It was shown by Taniyama [46] that only the torus knots $K=T(2,n)$ satisfy $u(K)=\lfloor(c(K)-1)/2\rfloor$. Abe, Hanaki, and Higa [2] showed that if $K$ is prime, then $u(K)=\lfloor(c(K)-2)/2\rfloor$ if and only if $K$ is either the knot $4_{1}$ (the figure-8 knot) or $K$ can be represented as a positive or negative 3-braid. They also showed that the only possible non-prime knots with $u(K)=\lfloor(c(K)-2)/2\rfloor$ are the knots T(2,n)#T(2,m)T(2,n)\#T(2,m) and $T(2,n)\#\overline{T(2,m)}$ for $n,m\geq 3$, and that the knots K=T(2,n)#T(2,m)K=T(2,n)\#T(2,m) do satisfy this equality. They conjectured that the knots $T(2,n)\#\overline{T(2,m)}$ also satisfy $u(K)=\lfloor(c(K)-2)/2\rfloor$, and showed that this is equivalent to the statement that $u(T(2,n)\#\overline{T(2,m)})=u(T(2,n))+u(\overline{T(2,m)})$. Our Corollary 1.3 shows that this last assertion is false, when $n,m\geq 7$.
This leaves the cases $T(2,3)\#\overline{T(2,m)}$ and $T(2,5)\#\overline{T(2,m)}$ for $m\geq 3$ and odd, to decide whether or not $u(K)=\lfloor(c(K)-2)/2\rfloor$ (that is, whether or not $u(T(2,3)\#\overline{T(2,m)})=1+\frac{m-1}{2}$ and $u(T(2,5)\#\overline{T(2,m)})=2+\frac{m-1}{2}$). Settling these last two cases would complete the characterization of the knots $K$ with $u(K)=\lfloor(c(K)-2)/2\rfloor$ .
@@ 3. How (and why) the main example was found
The process of finding the diagram for $7_{1}\#\overline{7_{1}}$ above and the discovery of its unexpectedly low unknotting number came as a consequence of another ongoing project of the authors: our search for an example to show that unknotting number is not invariant under mutation. This would answer Problem 1.69(C) in Kirby’s problem list [24], in the negative. Our approach to that problem included building a collection of knots which would have unknotting number 4 (if unknotting number were additive), but which would have comparatively small signature. The inequality $|\sigma(K)|/2\leq u(K)$ [29] is useful in bounding unknotting number from below, but it can be difficult to find knots where the difference between these two quantities is known to be greater than 1 (which we needed for our search), although work of Stoimenow [44] is a significant outlier, which deserves to be more widely known. On the other hand, the facts that signature is additive under connected sum, and that the signature of the mirror image of a knot $K$ is the negative of the signature of $K$ (see, e.g., [29]), means that is it a routine matter to build connected sums where the unknotting number would be large, if unknotting number were additive, but where the signature is small. It was for this reason that we began to search for knots a few crossing changes away from connected sums of low crossing number knots with known unknotting number, in order to build a collection of new knots with (if unknotting number were additive) high enough unknotting number to be useful for our project.
@@
Our approach to finding the new knots was to adopt the process we had followed for two previous projects, where we settled the Bernhard-Jablan Conjecture (in the negative) [11] and where we completed the computation of the smooth 4-genus of the prime knots through 12 crossings [12]. We started with pairs of knots, one with known unknotting number 3 and the other with known unknotting number 4, took their connected sum, and then built diagrams for these sums for which crossing changes stood a good chance of building prime knots, with low enough (after simplification) crossing number that SnapPy could recognize them (that is, crossing number 15 or less). We chose unknotting numbers 3 and 4 so that we were starting with connected sums that were likely to have unknotting number 7, and then randomly changed three crossings so that the resulting knots conjecturally would have unknotting number (at least) 4. We built the knot diagrams by starting with the standard diagrams for the connected sums, randomly performing some crossing-number-increasing Reidemeister moves using SnapPy’s backtrack() command, and then converting them to braids using the braid_word() command. For this last step, SnapPy uses Vogel’s algorithm [47] to convert the knot diagram into a closed braid. We chose this approach because (a) SnapPy can quickly carry out both of these operations and (b) our experience with the previous two projects had shown us that working with braid representatives built in this way succeeds in building diagrams that are randomized enough to yield good results (that is, many different knot types) when we change crossings in the braid representations. Experience also showed that changing 3 crossings, when starting from a connected sum in this way, resulted in a good proportion of the knots that were generated being prime; 2 crossing changes was typically insufficient. Using braid words, changing crossings is very quick; it amounts to changing the signs of entries in the braid representation.
@@
Our work in [11] had provided upper bounds on the unknotting number of all of the prime knots through 15 crossings, what we called, in that paper, their weak BJ-unknotting number, that is, their unknotting number if the Bernhard-Jablan conjecture were true. We could have confidence that the vast majority of these computations represented the actual unknotting number of the knots, based on comparison with known values for lower crossing number knots. We compared the knots coming from the randomized procedure described above to our list of ‘weak BJ-unknotting number 4’ knots to create a group of what we called ‘conditionally unknotting number 4’ knots. Working with several million braids yielded 240 such 14- and 15-crossing knots. One of these was the knot $K14a18636$.
@@
We then began to reverse the search, using these 240 knots as initial knots to carry out crossing changes on their braid representatives, in the same manner in which the connected sum diagrams had been built. For the mutation project, our goal was to show that one of a collection of target knots had unknotting number 3, by showing that it was crossing-adjacent to a knot with unknotting number 4. These target knots came from a database of mutant knot pairs constructed by by Stoimenow [45], using our Bernhard-Jablan data to identify mutant pairs with distinct weak BJ-unknotting number. Prior such searches using knots with unknotting number known to be 4 had failed to make such a connection, and so our plan was to search among knots one crossing change away from these 240 knots with conditional unknotting number 4, instead. Connecting one of the 240 initial knots with one of these mutant ‘targets’ would have proved that either unknotting number is not invariant under mutation, or that the knot from among our 240 that made the connection would actually have unknotting number at most 3, and so, as argued in Section 2, unknotting number would not be additive under connected sum. (But, were such a thing to happen, we would not know which of the two possibilities was true!)
@@
At the start of this phase of the project it was not clear what this process would produce. The most likely prospect seemed to be that we would never find a connection between our 240 initial knots and our list of target knots, if only because both conjectures might be true. Our expectation, if the project succeeded, was that it would be because the target knot we connected with had unknotting number 3, and so unknotting number would not be mutation invariant. It came as quite a surprise then when, instead, this reverse search found a crossing change from one of the 240 knots that we had identified, namely $K14a18636$, to the knot $K15n81556$, which had a known, and unexpectedly low, weak BJ-unknotting number of 2.
@@
The first example we found started from a closed braid representation of the knot $7_{1}\#10_{139}$; three crossing changes, that our search had found, transformed this to a representative of the knot $K14a18636$. One of those three crossing changes transformed the braid into a braid representative for the knot $7_{1}\#\overline{7_{1}}$, and so this braid word contained two crossing changes to a representative of the knot $K14a18636$. Initially, the braid word that made this connection had length 119, and so represented a 119-crossing knot diagram. But several simplifications of the braid closure could, it turned out, be carried out. These mostly involved Reidmeister I moves ‘hidden’ by a sequence of Reidemeister II moves. Their discovery was aided by the program KnotPlot [42], to visualize long stretches of the braid and find the crossings to eliminate. This enabled the authors to greatly reduce the number of crossings needed (from 119 to 20), and the number of strands in the braid needed (from 14 to 5). This resulted in the diagram and braid word used in our discussion in Section 2.
@@ 4. The road forward
With the additivity of unknotting number being settled in the negative, the quest to build a more comprehensive understanding of unknotting number becomes more challenging. The fact that unknotting number does not behave as expected with respect to connected sum means that building a purely computational approach to determining unknotting numbers, as one could argue the other invariants mentioned in the introduction enjoy, seems further out of reach. In particular, we cannot rely on knowledge of unknotting numbers for prime knots to provide all of the information we need to compute the invariant more generally.
As noted in the proof of Corollary 1.4 in Section 1, since the knot group of a connected sum is insensitive to mirroring, but its unknotting number can change depending on the mirroring of the summands, this means that the knot group alone is not sufficient to detect the unknotting number of knots. On the other hand, the knot group $\pi(K)=\pi_{1}(S^{3}\setminus K)$ together with the knot’s peripheral structure (the conjugacy class of the image of $\pi_{1}(\partial N(K))$ in $\pi(K)$ under the inclusion-induced homomorphism) does determine the knot $K$ [48, Corollary 6.5], and therefore (in the abstract) determines its unknotting number. Any general method to compute the unknotting number of a knot from data coming from the knot group would therefore need to take into account this peripheral structure. One approach might be to focus on the 2-generalized knot group $G_{2}(K)$, which is built from a presentation of the knot group and which is known [30] to determine the knot, up to mirroring. It therefore, in the abstract, determines the unknotting number, as well.
@@
How does $G_{2}(K)$ behave under connected sum? Can $u(K)$ be computed from $G_{2}(K)$?
Knowing that additivity of unknotting number can in fact fail makes it more interesting to discover when it does not fail. That is, establishing that $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for families of knots has increased relevance and importance, since we can no longer expect this to hold in general. As mentioned in Section 1, there are families $\mathcal{K}$} of knots for which additivity holds for pairs of knots chosen within the family [25],[5]; further classes would help give us a better picture of where we might look for more examples of non-additivity. This in turn might lead to new shortcuts, which would have relevance to the computation of unknotting number more generally.
Since additivity of unknotting number does not hold in general, we can also explore how badly additivity can fail. We can leverage the examples above to show that the difference u(K#L)u(K)u(L)u(K\#L)-u(K)-u(L) can be arbitrarily large, by, for example, using $K=\#_{n}7_{1}$ and $L=\#_{n}\overline{7_{1}}$, the $n$-fold connected sums of these knots. We have $u(K)=u(L)=3n$ by signature considerations, but by rearranging terms we have $K\#L=\#_{n}(7_{1}\#\overline{7_{1}})$, with unknotting number at most $5n$. Any knot that the knot $7_{1}$ is gordian adjacent to can be used in same fashion, and so the knots described in Corollaries 1.5 and 2.1 can be used to build infinitely many examples. This gives Corollary 1.6. $\square$
It is an interesting question to ask to what extent such behavior also occurs for prime knots $K$ and $L$. Sebastian Baader [8] has pointed out that the pretzel knot $L=P(7,7,\ldots,7,2)$ with $n$ 7’s, which is prime, can be transformed to the knot $_{n}7_{1}$ by a single crossing change in the last collection of half-twists. This implies that $u(L\#\overline{L})\leq 5n+2$; make the two needed crossing changes to reach $_{n}(7_{1}\#\overline{7_{1}}$). But $u(L)=u(\overline{L})\geq 3n-1$, since anything smaller would allow us to unknot $_{n}7_{1}$ more quickly. The $7$’s in this construction can also be replaced with any sequence of odd numbers $\geq 7$, while preserving the large additivity gap, by Corollary 1.3. What other examples can we build?
Even more than this, is it, perhaps, possible that the unknotting number of a connected sum can be lower than that of its summand?
@@
Does $u(K\#K^{\prime})\geq\max\{u(K),u(K^{\prime})\}$ for all knots $K$ and $K^{\prime}$ ?
The fact that we could bootstrap an initial, single, example of non-additivity to find infinitely may knots $K$ for which $u(K\#7_{1})<u(K)+u(7_{1})$ has shown that failure of additivity is far from an isolated phenomenon. This raises the question of whether or not every knot $K$ can be paired with a ‘poison’ knot $K^{\prime}$ so that $u(K\#K^{\prime})<u(K)+u(K^{\prime})$. So we ask:
@@
Does there exist a non-trivial knot $K$ for which $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for every knot $K^{\prime}$ ?
While we now know that $u(7_{1}\#\overline{7_{1}})<u(7_{1})+u(\overline{7_{1}})$, and we can build infinitely many examples of this phenomenon from this, all of our examples currently rely on a single sequence of $5$ crossing changes that unknot $7_{1}\#\overline{7_{1}}$ . We would certainly like to understand why this sequence occurred! What, if anything, is special about the knot $7_{1}$ that might make this happen? Understanding this unknotting sequence might help us understand how to find other methods to unknot knots more ‘efficiently’.
The intermediate knot $K_{1}=K14a18636$, which played a central role in establishing the non-additivity of unknotting number, has, from our work in [11], a weak BJ-unknotting number of 4, yet the unknotting sequence above shows that it has unknotting number at most 3. So the knot $K_{1}$ is, in fact, a counterexample to the weak Bernhard-Jablan conjecture [11]. That is, no crossing change in its (unique) minimal crossing diagram lowers the weak BJ-unknotting number. This may point the way to building more examples of unexpectedly low unknotting number, by trying to find a BJ-unknotting counterexample in an unknotting sequence. We note that there are (at least) six other knots that the authors have identified with weak BJ-unknotting 4 and unknotting number $\leq 3$: they are the knots $K13n1669$, $K14a2644$, $K14a11300$, $K14a11565$, $K14a12383$, and $K14a13809$. Each is crossing-adjacent to a 15-crossing diagram for a knot $K$ with $u(K)\leq 2$ (the corresponding unknotting number (at most) 2 knots are, respectively, $K15n9318$, $K15n9318$, $K15n35674$, $K15n82797$, $K15n4866$, and $K15n129095$), and so each of these knots has unknotting number at most $3$. These could, like $K14a18636$, be useful targets for building other unexpectedly short unknotting sequences for connected sums.
Finally, we end with the question:
Question 4.4.
What is the actual unknotting number of $7_{1}\#\overline{7_{1}}$ ?